# Efficient Test Data Generation for MC/DC with OCL and Search

### v1

@@ 1 Introduction
Aircraft avionics are safety-critical systems that are required to comply with international safety standards. DO-178C [1] is the prevailing standard for the avionics of aircraft systems. A large part of DO-178C (around 70%) focuses on verification and validation activities. Testing forms a significant portion of the overall activities required to comply with the standard. According to DO-178C, an important consideration for the highest critical level avionics applications is to test these applications based on the Modified Condition/Decision Coverage (MC/DC) criterion. The MC/DC criterion subsumes different coverage criteria, e.g., statement coverage, branch coverage, and multiple condition coverage. It is considered as one of the most rigorous coverage criteria. The empirical results of previous studies [110, 104, 37, 2] indicate that it is more effective in terms of fault detection as compared to other coverage criteria.
@@
The research work presented in this paper was conducted in collaboration with a large avionics systems industry as a part of the project for the National Center of Robotics and Automation (NCRA). Their main goal is to comply with international safety standards, i.e., DO-178C while testing avionics systems of manned and unmanned aircraft. International safety standards like DO-178C recommend coverage of requirements at both system and source code levels using stringent coverage criteria, e.g., MC/DC. OCL constraints specified on models at the system level are ultimately converted into conditions at the source code level [54]. Therefore, generating test data from OCL constraints targeting MC/DC coverage enables compliance with DO-178C test obligations, which is an important consideration of the avionics industry. The available techniques targeting OCL-based MC/DC test data generation highlight the need for optimizing resource and time [54]. Our goal, in this context, is to devise an efficient strategy for search-based MC/DC test data generation using OCL constraints specified on models. In this paper, we propose an effective way to automate MC/DC test data generation during model-based testing. For this purpose, we develop a strategy that utilizes case-based reasoning (CBR) [64] and range reduction heuristics designed to solve MC/DC-tailored OCL constraints. We performed an empirical study to compare our proposed strategy for MC/DC test data generation using CBR, range reduction, both CBR and range reduction, with an original search algorithm, and random search. Additionally, we conducted an experiment to compare our proposed CBR and range reduction methods with two constraint solvers, namely UMLtoCSP [19] and PLEDGE [95]. We performed a large-scale experiment using 129 OCL constraints from six case studies of varying nature and sizes. Among the six, one represents a Ground Control Station that is used to remotely control Unmanned Aerial Vehicles. The other two case studies, EU-Rental [35] and the Royal and Loyal [100, 17] are widely recognized in OCL literature as academic benchmarks [59]. The fourth case study represents the UML meta-model of State Machines [75]. The fifth case study is an artificial (Seat Belt System) that we developed inspired by various industrial case studies that are published in the literature [4, 6]. The sixth case study is industrial which represents a Student Registration System (SRS). We evaluated our proposed strategy using the search algorithm, Alternating Variable Method (AVM). The proposed strategy with CBR is termed as AVMc, and with range reduction, it is termed as AVMr. We compared our strategy with the original AVM search, which is termed AVMo. We used Random Search (RS) as a baseline for comparison. We also analyzed the effect of the combination of CBR and range reduction (AVMrc i.e., AVM with range reduction and CBR approach) and compared it with AVMc and AVMr. The experimental results showed that AVMo, AVMc, AVMr, and AVMrc outperform RS. The AVMc and AVMr performed significantly better when compared with AVMo in terms of the ability to solve OCL constraints. The results of the comparison with AVMrc indicated that the combination of CBR and range reduction (AVMrc) outperforms AVMr, AVMc, and AVMo. Based on the results, the combined strategy (AVMrc) is considered the best option for solving MC/DC constraints. Moreover, the results of the comparison with UMLtoCSP and PLEDGE showed that AVMrc outperformed by solving a higher percentage of MC/DC constraints. However, in terms of time efficiency, UMLtoCSP and PLEDGE demonstrated better performance.
This paper is an extension of our previous conference paper published in the Proceedings of 11th International Symposium on Search-Based Software Engineering [87]. The specific contributions of this paper in addition to the conference version are given below.
  • We introduced a range reduction method into the MC/DC test data generation strategy to efficiently solve MC/DC-tailored OCL constraints.
  • We significantly extended empirical evaluation in the following ways.
    • We added two new case studies (comprising 45 OCL constraints) and one new research question (RQ)
    • We extended the experiment reported in the conference paper (evaluation with CBR) for the two new case studies and additional RQ.
    • We evaluated the range reduction method using six case studies, 129 OCL constraints, and two RQs.
    • As a part of the new RQ, we analyzed our strategy with a combination of CBR and range reduction methods for all case studies.
  • We conducted an empirical evaluation comparing our proposed CBR and range reduction methods with two existing constraint solvers, i.e., UMLtoCSP and PLEDGE.
@@
The remaining paper is structured as follows. Section 2 provides the background of model-based testing, MC/DC, search-based software engineering, OCL constraints solving, and case-based reasoning. Section 3 presents the running example. Section 4 describes the proposed strategy to generate test data for MC/DC using CBR and range reduction. Section 5 presents the empirical evaluation of the proposed strategy. Section 6 discusses the works related to our paper. Finally, Section 7 concludes the paper.
@@ 2.1 Model-based Testing
Model-based testing (MBT) supports testing automation using an abstract representation of the system under test (SUT) in the form of models [97]. The process of MBT consists of five steps. In the first step, the specifications of the SUT are modeled using a modeling language, such as Unified Modeling Language (UML) [75]. To model various aspects of the system, UML provides several modeling artifacts for different purposes. The modeling artifacts provided by UML are broadly categorized as structural models (e.g., a class diagram and a profile diagram) and behavioral models (e.g., a state machine and an activity diagram). The models developed in UML are augmented using a constraint specification language, i.e., Object Constraint Language (OCL) [74]. OCL is a widely used textual language for specifying constraints on UML models [59, 90]. Different UML models support the automation of several testing activities. For example, the UML state machine can be used to generate test sequences [57, 88]. The second step is to select the coverage criteria for the test case selection. In the third step, the test case specification is defined based on the test selection criteria. In the fourth step, the test cases are generated using the model created in the first step, and the test case specification that is defined in the third step. Finally, in the fifth step, the executable test scripts for the SUT are created to execute the test cases generated in the previous step.
@@ 2.2 Modified Condition/Decision Coverage (MC/DC)
[…]
@@ 2.3 Search-based Software Engineering (SBSE)
In SBSE, a problem from the software engineering domain is formulated as a search problem, and then different search algorithms are used to solve that problem. A search problem formulation process typically consists of three steps [49]. The first step is to create a representation of the solution. The second step is to define a fitness function to evaluate the candidate solution during the search process. Finally, the third step is to use operators (e.g., cross-over) to evolve the candidate solution. The use of operators depends on the type of search algorithm (i.e., global or local). The global search algorithms (e.g., genetic algorithms) use the whole population (with more than one individual) to search. The local search algorithms (e.g., hill climbing) start the search process using one individual. For global search algorithms like genetic algorithms, cross-over, and mutation operators are used to generate offspring from two individuals. For local search algorithms like hill climbing, a mutation operator is needed to evolve an individual.
@@ 2.4 OCL Constraint Solving
The process of solving an OCL constraint involves generating values for the attributes that can satisfy the overall constraint. To generate values that can be used as test data, Ali et al. [4], proposed a search-based approach. The approach takes a UML class diagram and OCL constraints as inputs, uses a search algorithm to solve the constraints, and generates test data in the form of an object diagram as output. For this purpose, a set of heuristics was proposed to guide the search algorithms. A heuristic to calculate branch distance tells ‘how far’ input values are from satisfying the constraint. Based on the fitness value, a search algorithm iteratively improves input values to satisfy the given OCL constraint. For example, consider that we want to solve a constraint C: context ClassX inv: self.x ¿ 5 using the AVM algorithm. Suppose an initial random solution for self.x ¿ 15 is $x=0$. At the start, AVM evaluates this solution to analyze fitness. Since this solution is far from satisfying self.x ¿ 15, AVM will increment ($x=1$) and decrement ($x=-1$) the solution value. As the incremented value ($x=1$) is close to satisfying self.x ¿ 15, AVM will gradually improve the value of ‘x’ toward better fitness (i.e. close to the solution) and continues until a solution is found or some stopping criteria are met. For example constraint C, suppose after some iterations, the value of ‘x’ is $x=31$ which satisfies self.x ¿ 15. Now the solution is found and the AVM search process will stop. Finally, the solution for constraint C is saved as an object diagram corresponding to the input class diagram.
@@ 3 Running Example
This section presents a running example containing a set of OCL constraints. For this purpose, we use an industrial case study representing a Ground Control Station (GCS). A GCS is used to remotely operate an Unmanned Aerial Vehicle (UAV). One of the main functionalities of a GCS is to provide mission planning for the UAV. Figure 1 shows an excerpt of the GCS class diagram with five classes. The main class is GCS which is associated with UAV and Mission classes. Each mission has one Route and several Waypoints. Listing 1 shows 10 OCL constraints on various classes modeled for GCS. Among these constraints, five constraints (i.e., C1, C2, C3, C7, and C8) are related to the UAV mission, and the remaining five constraints (i.e., C4, C5, C6, C9, and C10) are modeled for route optimization.
The constraint C1 demonstrates that for each mission defined using GCS, the UAV’s flight time or flight distance must not exceed the maximum limit. C2 specifies that each mission must have at least a minimum number of waypoints. C3 defines a limit on long missions (e.g., for monitoring or surveying) that mission must have 100 waypoints in addition to the minimum limit for waypoints. C4 defines pre-condition on route optimization operation in which route distance must be greater than the minimum distance. Similar to C4, C5 and C6 define pre-conditions with margins of 1000 and 1500 feet respectively. C8 specifies UAV flight distance range with a margin of 100 meters. C9 defines a pre-condition on route optimization operation in which route distance must be greater than maximum range and input distance values should be valid. Lastly, C10 specifies a pre-condition on route optimization operation in which route distance with a margin must be greater than the total distance, and input distance values must be valid.
@@
--The number of waypoints in a mission should be greater than the minimum waypoints’ limit
@@ 4 Strategy to Generate Test Data for MC/DC
Generating test data for MC/DC requires the generation of multiple solutions for an OCL constraint corresponding to the MC/DC combinations. For example, consider a constraint with the predicate p $\lor$ q, where p and q are the clauses. To satisfy this predicate, there are four possible combinations for both p and q, i.e., TT, TF, FT, and FF. The main objective of the MC/DC criterion is to select a subset of only those combinations that have a direct impact on the outcome of the whole predicate. For this example, the combinations according to the MC/DC criterion are TF, FT, and FF.
Figure 2 shows an overall view of the proposed strategy to generate test data for MC/DC. As an initial step, the strategy takes an OCL constraint as input and reformulates it to obtain a list of MC/DC-tailored OCL constraints. The process of constraint reformulation involves handling the negation of logical expressions and the negation of OCL collections. The next step is to solve MC/DC-tailored OCL constraints using a search-based constraint solver. For this purpose, we propose two methods (as highlighted in Figure 2). The first method applies the concept of case-based reasoning (CBR) to optimize the process of constraint solving for the MC/DC-based constraints. The idea of CBR is to maintain a repository for solutions to reuse the previous solution while solving the current problem instead of solving it from scratch. Whenever a new MC/DC constraint is being solved, the repository is examined for the closest previous solutions. If no previous solution exists, the constraint solver generates a random solution and starts the search. If one or more closest previous solutions exist, the fitness of each previous and random solution is checked. This results in the closest solution that can be used as an initial seed of the search algorithm. The second method reduces the range of primitive attributes in the MC/DC constraints and generates random solutions within reduced ranges. These solutions are used by a search algorithm as a starting point to systematically change values during the search process. The aim is to minimize overall search steps which leads to further improvement in solving MC/DC constraints. Both methods can be used individually and collectively. At the end of the search process, the solution generated for the MC/DC constraint is stored in the repository for future reuse. In the following, we elaborate on these methods.
@@ 4.1.1 Handling Negation for OCL
4.1.2 Logical Operations Reformulation
To obtain the exact truth values that represent MC/DC combinations derived from the pair table, only negating various clauses is insufficient. The reason is that the heuristics (presented in [6]) for logical operations such as or, xor, and implies are developed to improve the search by solving the ‘easiest’ clause first. For example, in the case of predicate “A or B”, the heuristic to calculate the branch distance (as mentioned in [6]) is: If (d(A) <<<= d(B)) then solve A otherwise solve B clause. If the MC/DC combination requires the negation of clause B, it is possible that the negation of B is hard to solve as compared to clause A. In this case, the existing heuristics lead towards solving clause A. Therefore, it is not certain that the generated test data reflects the negation of B. The same is the case with xor and implies heuristics. To handle this, our strategy performs a second pass to reformulate the logical operations so that the generated test data directly reflects the MC/DC combination.

### v2

@@ 1 Introduction
Aircraft avionics are safety-critical systems that are required to comply with international safety standards. DO-178C [1] is the prevailing standard for the avionics of aircraft systems. A large part of DO-178C (around 70%) focuses on verification and validation activities. Testing forms a significant portion of the overall activities required to comply with the standard. According to DO-178C, an important consideration for the highest critical level avionics applications is to test these applications based on the Modified Condition/Decision Coverage (MC/DC) criterion. The MC/DC criterion subsumes different coverage criteria, e.g., statement coverage, branch coverage, and multiple condition coverage. It is considered one of the most rigorous coverage criteria. The empirical results of previous studies [110, 104, 37, 2] indicate that it is more effective in terms of fault detection as compared to other coverage criteria.
@@
The research work presented in this paper was conducted in collaboration with a large avionics systems industry as a part of the project for the National Center of Robotics and Automation (NCRA). Their main goal is to comply with international safety standards, i.e., DO-178C while testing avionics systems of manned and unmanned aircraft. International safety standards like DO-178C recommend coverage of requirements at both system and source code levels using stringent coverage criteria, e.g., MC/DC. OCL constraints specified on models at the system level are ultimately converted into conditions at the source code level [54]. Therefore, generating test data from OCL constraints targeting MC/DC coverage enables compliance with DO-178C test obligations, which is an important consideration of the avionics industry. The available techniques targeting OCL-based MC/DC test data generation highlight the need for optimizing resource and time [54]. Our goal, in this context, is to devise an efficient strategy for search-based MC/DC test data generation using OCL constraints specified on models.
In this paper, we propose an effective way to automate MC/DC test data generation during model-based testing. For this purpose, we develop a strategy that utilizes case-based reasoning (CBR) [64] and range reduction heuristics designed to solve MC/DC-tailored OCL constraints. We performed an empirical study to compare our proposed strategy for MC/DC test data generation using CBR, range reduction, both CBR and range reduction, with an original search algorithm, and random search. Additionally, we conducted an experiment to compare our proposed CBR and range reduction methods with two constraint solvers, namely UMLtoCSP [19] and PLEDGE [95]. We performed a large-scale experiment using 129 OCL constraints from six case studies of varying nature and sizes. Among the six, one represents a Ground Control Station that is used to remotely control Unmanned Aerial Vehicles. The other two case studies, EU-Rental [35] and the Royal and Loyal [100, 17] are widely recognized in OCL literature as academic benchmarks [59]. The fourth case study represents the UML meta-model of State Machines [75]. The fifth case study is an artificial (Seat Belt System) that we developed inspired by various industrial case studies that are published in the literature [4, 6]. The sixth case study is industrial which represents a Student Registration System.
We evaluated our proposed strategy using the search algorithm, Alternating Variable Method (AVM). The proposed strategy with CBR is termed AVMc, and with range reduction, it is termed AVMr. We compared our strategy with the original AVM search, which is termed AVMo. We used Random Search (RS) as a baseline for comparison. We also analyzed the effect of the combination of CBR and range reduction (AVMrc, i.e., AVM with range reduction and CBR approach) and compared it with AVMc and AVMr. The experimental results showed that AVMo, AVMc, AVMr, and AVMrc outperform RS. The AVMc and AVMr performed significantly better when compared with AVMo in terms of the ability to solve OCL constraints. The results of the comparison with AVMrc indicated that the combination of CBR and range reduction (AVMrc) outperforms AVMr, AVMc, and AVMo. Based on the results, the combined strategy (AVMrc) is considered the best option for solving MC/DC constraints. Moreover, the results of the comparison with UMLtoCSP and PLEDGE showed that AVMrc outperformed by solving a higher percentage of MC/DC constraints. However, in terms of time efficiency, UMLtoCSP and PLEDGE demonstrated better performance.
This paper is an extension of our previous conference paper published in the Proceedings of 11th International Symposium on Search-Based Software Engineering [87]. The specific contributions of this paper in addition to the conference version are as follows. We introduced a range reduction method into the MC/DC test data generation strategy to efficiently solve MC/DC-tailored OCL constraints. We significantly extended empirical evaluation in several ways. First, we incorporated two new case studies comprising 45 OCL constraints. Second, we added two new research questions (RQs). One of the RQs analyzes our strategy with a combination of CBR and range reduction methods for all case studies. Another RQ compares our proposed CBR and range reduction methods with two existing constraint solvers. Third, we extended the experiment reported in the conference paper (evaluation with CBR) for the two new case studies and two new RQs. Fourth, we evaluated our strategy with the range reduction method using six case studies, 129 OCL constraints, and all RQs. Lastly, we empirically evaluated our proposed methods with two existing constraint-solving approaches, i.e., UMLtoCSP and PLEDGE.
@@
The remaining paper is structured as follows. Section 2 provides the background of model-based testing, MC/DC, search-based software engineering, OCL constraints solving, and case-based reasoning. Section 3 presents a running example. Section 4 describes our proposed strategy to generate test data for MC/DC using CBR and range reduction. Section 5 presents the empirical evaluation of our proposed strategy. Section 6 discusses the works related to our paper. Finally, Section 7 concludes the paper.
@@ 2.1 Model-based Testing
Model-based testing (MBT) supports testing automation using an abstract representation of the system under test (SUT) in the form of models [97]. The process of MBT consists of five steps. In the first step, the specifications of the SUT are modeled using a modeling language, such as Unified Modeling Language (UML) [75]. To model various aspects of the system, UML provides several modeling artifacts for different purposes. The modeling artifacts provided by UML are broadly categorized as structural models (e.g., a class diagram and a profile diagram) and behavioral models (e.g., a state machine and an activity diagram). The models developed in UML are augmented using a constraint specification language, i.e., Object Constraint Language (OCL) [74]. OCL is a widely used textual language for specifying constraints on UML models [59]. Different UML models support the automation of several testing activities. For example, the UML state machine can be used to generate test sequences [57, 88]. The second step is to select the coverage criteria for the test case selection. In the third step, the test case specification is defined based on the test selection criteria. In the fourth step, the test cases are generated using the model created in the first step, and the test case specification that is defined in the third step. Finally, in the fifth step, the executable test scripts for the SUT are created to execute the test cases generated in the previous step.
@@ 2.2 Modified Condition/Decision Coverage (MC/DC)
MC/DC is a type of structural coverage criterion used in software testing. This criterion subsumes other structural coverage criteria such as statement coverage, decision coverage, condition coverage, and condition/decision coverage [25]. Therefore, it is considered a stronger coverage criterion [39]. It is a widely used criterion in the industry mainly in the avionics domain for testing safety and mission-critical applications. It is also one of the main requirements of various international safety standards such as DO-178C [1]. Chilenski and Miller [25] defined MC/DC as: “Modified condition/decision coverage is a structural coverage criterion requiring that each condition within a decision is shown by execution to independently and correctly affect the outcome of the decision.”. For example, for a decision ‘$X$ $and$ $Y$’, the possible combinations of each condition are TT, TF, FT, and FF. Out of these combinations, only TT, TF, and FT affect the decision, i.e., by changing $X/Y$ from T to F, or inversely, the decision outcomes change. These three combinations satisfy the MC/DC criterion and can be selected for testing. As the number of conditions in a decision increases, the total number of combinations increases by 2n. This results in many combinations to test, which is infeasible for large-scale applications, e.g., safety-critical systems. MC/DC was designed to enable rigorous testing of safety-critical systems and comply with international safety standards [25].
@@ 2.3 Search-based Software Engineering (SBSE)
In SBSE, a problem from the software engineering domain is formulated as a search problem, and then different search algorithms are used to solve that problem. A search problem formulation process typically consists of three steps [49]. The first step is to create a representation of the solution. The second step is to define a fitness function to evaluate the candidate solution during the search process. Finally, the third step is to use operators (e.g., crossover) to evolve the candidate solution. The use of operators depends on the type of search algorithm (i.e., global or local). The global search algorithms (e.g., genetic algorithms) use the whole population (with more than one individual) to search. The local search algorithms (e.g., hill climbing) start the search process using one individual. For global search algorithms like genetic algorithms, crossover, and mutation operators are used to generate offspring from two individuals. For local search algorithms like hill climbing, a mutation operator is needed to evolve an individual.
@@ 2.4 OCL Constraint Solving
The process of solving an OCL constraint involves generating values for the attributes that can satisfy the overall constraint. To generate values that can be used as test data, Ali et al. [4], proposed a search-based approach. The approach takes a UML class diagram and OCL constraints as inputs, uses a search algorithm to solve the constraints, and generates test data in the form of an object diagram as output. For this purpose, a set of heuristics was proposed to guide the search algorithms. A heuristic to calculate branch distance tells ‘how far’ input values are from satisfying the constraint. Based on the fitness value, a search algorithm iteratively improves input values to satisfy the given OCL constraint. For example, consider that we want to solve a constraint C: context ClassX inv: self.x ¿ 15 using the AVM algorithm. Suppose an initial random solution for self.x ¿ 15 is $x=0$. At the start, AVM evaluates this solution to analyze fitness. Since this solution is far from satisfying self.x ¿ 15, AVM will increment ($x=1$) and decrement ($x=-1$) the solution value. As the incremented value ($x=1$) is close to satisfying self.x ¿ 15, AVM will gradually improve the value of ‘x’ toward better fitness (i.e., close to the solution) and continue until a solution is found or some stopping criteria are met. For example constraint C, suppose after some iterations, the value of ‘x’ is $x=31$ which satisfies self.x ¿ 15. Now the solution is found and the AVM search process will stop. Finally, the solution for constraint C is saved as an object diagram corresponding to the input class diagram.
@@ 3 Running Example
This section presents a running example containing a set of OCL constraints. For this purpose, we use an industrial case study representing a Ground Control Station (GCS). A GCS is used to remotely operate an Unmanned Aerial Vehicle (UAV). One of the main functionalities of a GCS is to provide mission planning for the UAV. Figure 1 shows an excerpt of the GCS class diagram with five classes. The main class is GCS which is associated with UAV and Mission classes. Each mission has one Route and multiple Waypoints. Listing 1 shows 10 OCL constraints on various classes modeled for GCS. Among these constraints, five constraints (i.e., C1, C2, C3, C7, and C8) are related to the UAV mission, and the remaining five constraints (i.e., C4, C5, C6, C9, and C10) are modeled for route optimization.
The constraint C1 demonstrates that for each mission defined using GCS, the UAV’s flight time or flight distance must not exceed the maximum limit. C2 specifies that the number of waypoints in a mission must exceed the minimum threshold for the number of waypoints. C3 defines a limit on long missions (e.g., for monitoring or surveying) that mission must have 100 waypoints in addition to the minimum limit for waypoints. C4 defines pre-condition on route optimization operation in which route distance must be greater than the minimum distance. Similar to C4, C5 and C6 define pre-conditions with margins of 1000 and 1500 feet respectively. C8 specifies UAV flight distance range with a margin of 100 meters. C9 defines a pre-condition on route optimization operation in which route distance must be greater than maximum range and input distance values should be valid. Lastly, C10 specifies a pre-condition on route optimization operation in which route distance with a margin must be greater than the total distance, and input distance values must be valid.
@@
--The number of waypoints in a mission should be greater than the minimum number of waypoints
@@ 4 Strategy to Generate Test Data for MC/DC
Generating test data for MC/DC requires the generation of multiple solutions for an OCL constraint corresponding to the MC/DC combinations. For example, consider a constraint with a predicate p $\lor$ q, where p and q are the clauses. To satisfy this predicate, there are four possible combinations for both p and q, i.e., TT, TF, FT, and FF. For this example, the combinations according to the MC/DC criterion are TF, FT, and FF.
Figure 2 shows an overall view of the proposed strategy to generate test data for MC/DC. As an initial step, the strategy takes an OCL constraint as input and reformulates it to obtain a list of MC/DC-tailored OCL constraints. The process of constraint reformulation involves handling the negation of logical expressions and the negation of OCL collections. The next step is to solve MC/DC-tailored OCL constraints using a search-based constraint solver. For this purpose, we propose two methods (as highlighted in Figure 2). The first method applies the concept of case-based reasoning (CBR) to optimize the process of constraint solving for the MC/DC-based constraints. The idea of CBR is to maintain a repository for solutions to reuse the previous solution while solving the current problem instead of solving it from scratch. Whenever a new MC/DC constraint is being solved, the repository is examined for the closest previous solutions. If no previous solution exists, the constraint solver generates a random solution and starts the search. If one or more closest previous solutions exist, the fitness of each previous and random solution is checked. This results in the closest solution that can be used as an initial seed of a search algorithm. The second method reduces the range of primitive attributes in the MC/DC constraints and generates random solutions within reduced ranges. These solutions are used by a search algorithm as a starting point to systematically change values during the search process. The aim is to minimize overall search steps which leads to further improvement in solving MC/DC constraints. Both methods can be used individually and collectively. At the end of the search process, the solution generated for the MC/DC constraint is stored in the repository for future reuse. In the following, we elaborate on the steps of our strategy.
@@ 4.1.1 Handling Negation for OCL
4.1.2 Logical Operations Reformulation
To obtain the exact truth values that represent MC/DC combinations derived from the pair table, only negating various clauses is insufficient. The reason is that the heuristics (presented by Ali et al. [6]) for logical operations such as or, xor, and implies are developed to improve the search by solving the ‘easiest’ clause first. For example, in the case of predicate “A or B”, the heuristic to calculate the branch distance [6] is: If (d(A) <<<= d(B)) then solve A otherwise solve B clause. If the MC/DC combination requires the negation of clause B, it is possible that the negation of B is hard to solve as compared to clause A. In this case, the existing heuristics lead towards solving clause A. Therefore, it is not certain that the generated test data reflects the negation of B. The same is the case with xor and implies heuristics. To handle this, our strategy performs a second pass to reformulate the logical operations so that the generated test data directly reflects the MC/DC combination.